Thursday, 27 August

Plenary Lectures

9:00 – 10:00  Auditorium Maximum H 105

Marcelo Viana, IMPA, Rio de Janeiro, Brazil Dynamics: A Probabilistic and Geometric Perspective (I: p.25)

Chair: Jacob Palis, IMPA, Rio de Janeiro, Brazil

10:15 – 11:15  Auditorium Maximum H 105

Stéphane Mallat, École Polytechnique, CMAP, Palaiseau, France and Courant Institute, New York University, New York, USA Applied Mathematics Meet Signal Processing (I: p.13)

Chair: David Mumford, Brown University, Providence, USA

11:45 – 12:45  Auditorium Maximum H 105

Peter Sarnak, Princeton University, Princeton, USA Zeta and L-Functions (I: p.19)

Chair: John Coates, University of Cambridge, Cambridge, UK

14:00 – 15:00  Auditorium Maximum H 105

Persi Diaconis, Mathematics and ORIE, Cornell University, Ithaca, USA and Departments of Mathematics and Statistics, Stanford University, Stanford, USA From Shuffling Cards to Walking Around the Building (I: p.6)

Chair: David Siegmund, Stanford University, Stanford, USA

Closing Ceremony

15:00 – 16:00  Auditorium Maximum H 105

Program to be announced

ICM’98 – Program  1